the wattage rating W (in watt) of an appliance varies jointly with the square of the current I ( in amperes) and the resistance R ( in ohms). if the wattage is 6 watts when the current is 0.2 ampere and the resistance is 150 ohms, find the wattage when the ccurrent is 0.3 ampere and the resistance is 300 ohms
Answer by mananth(16949)

the wattage rating W (in watt) of an appliance varies jointly with the square of the current I ( in amperes) and the resistance R ( in ohms). if the wattage is 6 watts when the current is 0.2 ampere and the resistance is 150 ohms, find the wattage when the ccurrent is 0.3 ampere and the resistance is 300 ohms
...
Walphai^2r
W= ki^2r
if the wattage is 6 watts when the current is 0.2 ampere and the resistance is 150 ohms
6=k*(0.2)^2*150
k=6/(0.2)^2*150
k= 1
..
ccurrent is 0.3 ampere and the resistance is 300 ohms
W= 1*(0.3)^2*300
w= 27
...
m.ananth@hotmail.ca